Assume MM world with corporate tax ?c=40%. Risk free rate rf = 4%, market risk premium

= 10%. If a firm is unlevered, equity beta is 1.6. Assume that he firm issues debt and repurchases

equity with the proceeds and that the new D/E = 0.25 and return on debt rD = 6%. Find new

WACC.
Answer #1
Levered Beta = Unlevered Beta x (1 + ((1 – Tax Rate) x (Debt/Equity)))
levered beta = 1.6*(1+((1-0.4)*(0.25)))
levered beta = 1.84
D/A = D/(E+D)
D/A = 0.25/(1+0.25)
=0.2
Weight of equity = 1-D/A
Weight of equity = 1-0.2
W(E)=0.8
Weight of debt = D/A
Weight of debt = 0.2
W(D)=0.2
Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (Market risk premium)
Cost of equity% = 4 + 1.84 * (10)
Cost of equity% = 22.4
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 6*(1-0.4)
= 3.6
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=3.6*0.2+22.4*0.8
WACC =18.64%